# Compression of quantum shallow-circuit states

### v1

@@ Abstract
Shallow quantum circuits feature not only computational advantage over their classical counterparts but also cutting-edge applications. Storing quantum information generated by shallow circuits is a fundamental question of both theoretical and practical importance that remained largely unexplored. In this work, we show that $N$ copies of an unknown $n$-qubit state generated by a fixed-depth circuit can be compressed into a hybrid memory of $O(n\log_{2}N)$ (qu)bits, which achieves the optimal scaling of memory cost. Our work shows that the computational complexity of resources can significantly impact the rate of quantum information processing, offering a unique and unified view of quantum Shannon theory and quantum computing in the NISQ era.
I Introduction
I.1 Background
@@
Shallow quantum circuits are a focus of recent research, for they are arguably the most accessible resources with genuine quantum features and advantages. At the fundamental level, shallow quantum circuits with constant depth have been shown to be hard to simulate classically (unless BQP $\subseteq$ AM) [1], and they outperform their classical counterparts in certain computational tasks [2, 3]. In practice, variational shallow circuits [4, 5, 6, 7] will remain a core ingredient of quantum algorithms in the noisy and intermediate-scale quantum (NISQ) era [8]. Efficient methods of learning shallow and bounded-complexity quantum circuits have recently been proposed [9, 10, 11].
@@
Here we ask a fundamental question: Given $N$ copies of an unknown $n$-qubit state and the promise that it is generated by a shallow circuit, is there a faithful compression protocol that encodes the $N$-copy state into a memory of fewer (qu)bits and then decodes it up to an error vanishing at $N\to\infty$? Processing quantum states in the many-copy form is important for extracting, storing, and distributing quantum information. Tasks where many-copy states serve as a fundamental resource, to list a few, include quantum metrology [12, 13, 14], quantum state tomography [15, 16] and shadow tomography [17, 18], quantum cloning [19, 20, 21, 22], and quantum hypothesis testing [23, 24, 25, 26]. Quantum algorithms, such as quantum principle component analysis [27], may also require states in the many-copy form. As such, compression of quantum states in the many-copy form is a basic and crucial protocol required for their storage and transmission. In the literature, compression of many-copy states was first studied for the simple case of pure qubits by Plesch and Bužek [28], experimentally demonstrated in Ref. [29], and later generalized in a series of works to mixed qudits [30, 31, 32, 33]. However, regarding states generated by shallow quantum circuits, the existing results are not applicable, for they all assume the state to be in a fixed-dimension space. Here, instead, we consider states in a growing-dimension ($D=2^{n}$) space with complexity constraints. Therefore, studying the compression of shallow-circuit states not only requires better understanding of this important family of states but also demands new techniques of asymptotic quantum information processing.
I.2 Overview of main results
@@
Here $d_{\operatorname{Tr}}$ denotes the trace distance between quantum states. The encoder $\mathcal{E}_{N}$ and the decoder $\mathcal{D}_{N}$ are dependent on $N$ but are independent of the input state. The memory cost is characterized by the dimension of the Hilbert space spanned by $\{\mathcal{E}_{N}(\rho^{\otimes N})\}$. The goal of compression is to reduce the memory cost
$M:=\log_{2}\left|\mathsf{Supp}\left\{\mathcal{E}_{N}(\rho^{\otimes N})\right\}_{\rho\in\mathsf{S}}\right|,$ (2)
i.e., the number of (qu)bits required for storing $\mathcal{E}_{N}(\rho^{\otimes N})$, while respecting the faithfulness condition (1).
@@
Here, we are interested in the set of shallow-circuit states $\mathsf{S}_{{\rm sc}}$, which contains all $n$-qubit pure states that can be generated from $|0\rangle^{\otimes n}$ by circuits of depth no more than a constant $d$. As a proof-of-principle example, we focus on the most representative case of brickwork shallow circuits and consider the set of shallow-circuit states
$\mathsf{S}_{{\rm sc}}:=\left\{|\psi\rangle~{}:~{}|\psi\rangle=U_{\rm sc}|0\rangle^{\otimes n}\,\exists\,U_{{\rm sc}}\right\},$ (3)
where $U_{\rm sc}$ is a brickwork circuit with bounded depth $(\leq d)$ (as illustrated later in Figure 1).
Without compression, the memory cost of storing the input state equals $N\cdot n$ qubits. Our main contribution is to show that a faithful $N$-copy compression exists for $\mathsf{S}_{{\rm sc}}$, as long as $N$ grows at least as a polynomial of $n$ with a high enough degree (see Theorem 3 in Section V for details). The memory cost of the compression is linear in $n$ and logarithmic in $N$, i.e., $M=O(n\cdot\log_{2}N)$, achieving an exponential memory reduction in terms of $N$. Moreover, the memory does not have to be fully quantum. Instead, one may use a classical-quantum hybrid memory, where the ratio between the number of qubits and the number of classical bits decreases as $O(\log_{2}n/\log_{2}N)$. That is to say, when $N$ is large, the memory consists mainly of classical bits (while a fully classical memory doesn’t work; see Theorem 5 in Subsection V.6).
Following the main result, it is natural to ask if the memory cost can be further reduced. We prove (see Theorem 4 in Subsection V.5) that a memory of size $\Omega(n\cdot\log_{2}N)$ is required for keeping the compression faithful. In this sense, our compression protocol is optimal in the scaling of $n$ and $N$.
To establish the compression protocol, we develop novel tools for quantum information processing in the asymptotic regime of many copies, including a method of parameterizing shallow-circuit states in a small neighborhood with only $\mathsf{poly}(n)$ parameters (see Section III) and a correspondence between copies of a low-complexity state and a multi-mode coherent state (see Section IV). These tools can be further applied to other information processing tasks involving complexity-constrained quantum states.
I.3 Discussions
We have shown that $N$ copies of an $n$-qubit shallow-circuit state can be optimally compressed to $\Theta(n\cdot\log_{2}N)$ qubits. Intriguingly, the two key parameters $n$ (the number of qubits per copy) and $N$ (the number of copies) take distinct positions in the compression rate. We may give an interpretation to this phenomenon: $n$ is the parameter of informativeness, as it is proportional to the number of free parameters of a shallow-circuit state. On the other hand, $N$ is the parameter of accuracy, since $1/\sqrt{N}$ is the error scaling of tomography, i.e., of how well can we learn the information in the state. Our result shows that the $N$-copy state can be exponentially compressed only in the parameter of accuracy.
Besides memory efficiency, one may also be curious about the computational efficiency of shallow-circuit state compression. Unfortunately, the compression protocol in this work, despite being memory-efficient, is not computational efficient (see Subsection V.7). The main obstacle is that the protocol requires searching over a covering mesh of shallow-circuit states, whose cost is exponentially large (in $n$). It is noteworthy that this is also the key step of converting a part of the memory to classical bits. It is thus intriguing to conjecture that any protocol using a hybrid memory is computationally inefficient. On the other hand, there exist compression protocols using fully quantum memory [30, 31] that do not require searching, and there remains hope that these protocols could inspire a computationally efficient protocol for shallow-circuit states.
@@ I.3 Discussions
As we focused on the most fundamental case, there is plenty room for extension. For example, one may consider shallow-circuit states with a 2D structure, and the techniques developed here should apply. Moreover, the circuit depth $d$ is treated as a constant throughout this work, but from the derivation of results it can be seen that the compression will still be faithful when $d$ grows very slowly (e.g., $d\ll\log n$) with $n$. In particular, it would be interesting to cover pesudorandom quantum states [34], which are low-depth states processing approximate Haar-randomness and are thus of particular interest in quantum cryptography. At last, one may even take into account the effect of noise and consider the compression of noisy shallow-circuit states. While similar results are expected there, some techniques in this work do not immediately generalize to mixed states and require moderate adaptation.

### v3

@@ Abstract
Shallow quantum circuits feature not only computational advantages over their classical counterparts but also cutting-edge applications. Storing quantum information generated by shallow circuits is a fundamental question of both theoretical and practical importance that remained largely unexplored. In this work, we show that $N$ copies of an unknown $n$-qubit state generated by a fixed-depth circuit can be compressed into a hybrid memory of $O(n\log_{2}N)$ (qu)bits, which achieves the optimal scaling of memory cost. Our work shows that the computational complexity of resources can significantly impact the rate of quantum information processing, offering a unique and unified view of quantum Shannon theory and quantum computing.
@@
Introduction. Shallow quantum circuits are a focus of recent research, for they are arguably the most accessible resources with genuine quantum features and advantages. Fundamentally, shallow quantum circuits with constant depth are hard to simulate classically (unless BQP $\subseteq$ AM) [1], and they outperform their classical counterparts in certain computational tasks [2, 3]. Practically, variational shallow circuits [4, 5, 6, 7] will remain a core ingredient of quantum algorithms in the noisy and intermediate-scale quantum (NISQ) era [8]. Efficient methods of learning shallow and bounded-complexity quantum circuits have recently been proposed [9, 10, 11].
@@
Here we ask a fundamental question: Given $N$ copies of an unknown $n$-qubit state and the promise that it is generated by a shallow circuit, is there a faithful compression protocol that encodes the $N$-copy state into a memory of fewer (qu)bits and then decodes it up to an error vanishing at $N\to\infty$? Processing quantum states in the many-copy form is important for extracting, storing, and distributing quantum information. Tasks where many-copy states serve as a fundamental resource, to list a few, include quantum metrology [12, 13, 14], state tomography [15, 16] and shadow tomography [17, 18], quantum cloning [19, 20, 21, 22], and hypothesis testing [23, 24, 25, 26]. Quantum algorithms such as quantum principle component analysis [27] also require states in the many-copy form. As such, compression of many-copy quantum states is a basic and crucial protocol required for their storage and transmission. In the literature, compression of many-copy states was first studied for pure qubits [28], experimentally demonstrated in Ref. [29], and later generalized in a series of works to mixed qudits [30, 31, 32, 33]. However, regarding states generated by shallow quantum circuits, the existing results are not applicable, for they all assume the state to be in a fixed-dimension space. Here, instead, we consider states in a growing-dimension ($D=2^{n}$) space. Therefore, studying the compression of shallow-circuit states not only requires a better understanding of this important family of states but also demands new techniques of asymptotic quantum information processing.
Figure 1: We show that $N$ copies of an $n$-qubit state generated by an unknown brickwork shallow circuit can be compressed into a classical-quantum hybrid memory of size $O(n\log_{2}N)$ with vanishing error.
Without compression, storing the input state requires $N\cdot n$ qubits. Our main contribution, as illustrated in Figure 1, is to show that a faithful $N$-copy compression exists for brickwork shallow-circuit states, as long as $N$ grows at least as a polynomial of $n$ with a high enough degree. The memory cost of the compression is linear in $n$ and logarithmic in $N$, achieving the optimal rate with an exponential memory reduction in terms of $N$. Intriguingly, the memory does not have to be fully quantum. Instead, a classical-quantum hybrid memory suffices, where the ratio between the number of qubits and the number of classical bits decreases as $O(\log_{2}n/\log_{2}N)$. That is to say, when $N$ is large, the memory consists mainly of classical bits, while we also prove that a fully classical memory does not work. The compression is achieved with novel tools, which we develop for quantum information processing in the asymptotic regime of many copies. They can be further applied to other information-processing tasks involving complexity-constrained quantum states.
The setup. For a pure state $|\psi\rangle$, $\psi$ denotes the projector $|\psi\rangle\langle\psi|$. For a unitary $U$, the corresponding calligraphic letter $\mathcal{U}$ denotes the corresponding channel $\mathcal{U}(\cdot):=U(\cdot)U^{\dagger}$. The big-$O$, big-$\Omega$, and big-$\Theta$ notation  [^1] For two real-valued functions $f,g$, $f=O(g)$ if there exist $n_{0}$ and $c>0$ such that $|f(n)|\leq c\cdot g(n)$ for every $n\geq n_{0}$, $f=\Omega(g)$ if $g=O(f)$, and $f=\Theta(g)$ if both $f=O(g)$ and $f=\Omega(g)$. In addition, we denote by $f\ll g$ (which is also denoted by $f=o(g)$ in the literature) for non-negative functions $f$ and $g$ if $f(n)/g(n)$ vanishes in the limit of large $n$. will be used to discuss the asymptotic behavior of parameters. A function $f=f(n)$ is polynomial in $n$, denoted by $f\in\mathsf{poly}(n)$, if the exists a constant $k\geq 0$ such that $f=O(n^{k})$.
@@
Here $d_{\operatorname{Tr}}$ denotes the trace distance between quantum states. We focus on the blind setting [35, Chapter 10], where $\mathcal{E}_{N}$ and $\mathcal{D}_{N}$ are dependent on $N$ but are independent of the input state. The goal of compression is to reduce the memory cost $M:=\log_{2}\left|\mathsf{Supp}\left\{\mathcal{E}_{N}(\rho^{\otimes N})\right\}_{\rho\in\mathsf{S}}\right|$, i.e., the number of (qu)bits required for storing $\mathcal{E}_{N}(\rho^{\otimes N})$, while respecting the faithfulness condition (1).
@@
Here, we are interested in $n$-qubit pure states generated by circuits of depth no more than $d$. As a proof-of-principle example, we focus on the most representative case of (1D) brickwork shallow circuits and consider the set of shallow-circuit states $\mathsf{S}_{{\rm sc}}:=\left\{|\psi\rangle~{}:~{}|\psi\rangle=U_{\rm sc}|0\rangle^{\otimes n}\,\exists\,U_{{\rm sc}}\right\}$, where $U_{\rm sc}$ is a brickwork circuit with bounded depth $(\leq d)$ (as illustrated in Figure 1). We treat $n$ as the asymptotic parameters, $N$ as a growing function of $n$, and $d$ as a fixed parameter.
We begin with a sketch of the idea. First, by performing tomography on a negligible portion of the copies (while keeping most of the $N$ copies untouched), we localize the compression problem by fixing the unknown shallow-circuit state to a neighborhood (i.e., a region with radius vanishing in $N$). We then show that all shallow-circuit states in this neighborhood can be generated by a circuit consisting of $O(n)$ small rotations plus a fixed shallow circuit, thereby locally parameterizing this neighborhood by $O(n)$ parameters. We proceed to establish an approximate equivalence between copies of a state within this neighborhood and a (multi-mode) coherent state of light, which can be compressed by photon number truncation. More details are as follows.
Efficient local parameterization of shallow-circuit states. The first ingredient is a tool to pin the shallow-circuit state $|\psi_{{\rm sc}}\rangle$ down to a small local region parameterised by up to $\mathsf{poly}(n)$ real parameters. We call this task the local parameterization of quantum states, which is achieved by obtaining a crude estimate and then constructing a local model that incorporates the remaining uncertainty of the state. The first step requires shallow-circuit state tomography, e.g., [9]. The other step, however, is rather tricky as one has to parameterise the estimated neighborhood with as few parameters as possible.
We begin with general definitions. A unitary gate $U^{\epsilon}$ is called an $\epsilon$-rotation if $d_{\diamond}(\mathcal{U}^{\epsilon},\mathcal{I})\leq\epsilon$, where $d_{\diamond}$ denotes the diamond distance. An $\epsilon$-rotation can be expressed as a short-time unitary evolution $e^{-i2\epsilon H}$ for some Hermitian $H$ with $\|H\|_{\infty}\leq 1$ (see Appendix for the proof).
Definition 1.
Given a set $\mathsf{S}$ of $n$-qubit quantum states and a fixed quantum state $|\psi_{0}\rangle=U_{0}|0\rangle^{\otimes n}$, with the guarantee that $d_{\operatorname{Tr}}(\psi,\psi_{0})\leq\epsilon$ for every $|\psi\rangle\in\mathsf{S}$, the task of local parameterization of $\mathsf{S}$ is to identify a set $\mathsf{W}_{\rm loc}$ of $O(\epsilon)$-rotations, such that for every $|\psi\rangle\in\mathsf{S}$ there exists a $W\in\mathsf{W}_{\rm loc}$ with $|\psi\rangle=U_{0}W|0\rangle^{\otimes n}$. Moreover, the local parameterization is efficient if $\mathsf{W}_{\rm loc}$ can be parameterized with $\mathsf{poly}(n)$ local parameters, i.e., if there exists an on-to function from $\mathsf{E}^{f}$ to $\mathsf{W}_{\rm loc}$ with $\mathsf{E}\subset\mathbb{R}$ being an $O(\epsilon)-$ length interval and $f\in\mathsf{poly}(n)$.
It is quite easy to find a local parameterization: Applying $(U_{0})^{\dagger}$ on every $|\psi\rangle$, the similarity of $U_{0}^{\dagger}|\psi\rangle$ and $|0\rangle^{\otimes n}$ guarantees the existence of an $O(\epsilon)$-rotation. However, achieving an efficient parameterization is nontrivial. For an $n$-qubit state, the above naive strategy leaves the state in a neighborhood with exponentially many parameters. A hope is to improve parameter efficiency when the set of states under consideration has a certain structure. For instance, when $\mathsf{S}$ consists of states generated by locally acting an unknown quantum gate on a few qubits of a fixed $n$-qubit state $|\psi_{0}\rangle$. Suppose we know that all states in $\mathsf{S}$ are close to $|\psi_{0}\rangle$. It is quite natural to speculate that we could replace the few-qubit gate by an $O(\epsilon)$-rotation, acting on the same qubits, to achieve an efficient local parameterization (since the $O(\epsilon)$-rotation acts only on a few qubits, it can be modeled by fewer parameters). Somewhat surprisingly, this idea fails: There exists a bipartite pure state $|\psi\rangle_{AB}$ and a unitary $V_{B}$ acting on $B$ such that 1)1)1 ) $d_{\operatorname{Tr}}((\mathcal{I}_{A}\otimes\mathcal{V}_{B})(\psi_{AB}),\psi_{AB})\leq\epsilon$ and 2)2)2 ) there is no $O(\epsilon)$-rotation $V_{B}^{\epsilon}$ such that $(\mathcal{I}_{A}\otimes\mathcal{V}_{B})(\psi_{AB})=(\mathcal{I}_{A}\otimes\mathcal{V}_{B}^{\epsilon})(\psi_{AB})$ (see  [^2] One such example is $|\psi\rangle_{AB}=\sqrt{1-\epsilon}|0\rangle_{A}|00\rangle_{B}+\sqrt{\epsilon}|1\rangle_{A}\frac{|10\rangle_{B}+|20\rangle_{B}}{\sqrt{2}}$ (where $A$ is a qubit system and $B$ consists of two qutrit systems) and a gate $V_{B}$ that acts trivially on the second qutrit if the first qutrit is in $|0\rangle$ and maps the second qutrit from $|0\rangle$ to $|1\rangle$ ($|2\rangle$) conditioning on the first qutrit being $|1\rangle$ ($|2\rangle$). The output state is $\sqrt{1-\epsilon}|0\rangle_{A}|00\rangle_{B}+\sqrt{\epsilon}|1\rangle_{A}|B_{0}\rangle_{B}$ (with $|B_{0}\rangle:=\frac{|11\rangle+|22\rangle}{\sqrt{2}}$). If a desired $O(\epsilon)$-rotation $V^{\epsilon}$ on $B$ exists, it has to take the initial marginal state on $B$ to the marginal of the output state, i.e., $(1-\epsilon)|00\rangle\langle 00|+\epsilon\left(\frac{|10\rangle+|20\rangle}{\sqrt{2}}\right)\left(\frac{\langle 10|+\langle 20|}{\sqrt{2}}\right)\xrightarrow{V^{\epsilon}}(1-\epsilon)|00\rangle\langle 00|+\epsilon|B_{0}\rangle\langle B_{0}|$. Since a unitary cannot change the eigenvalues, we need $V^{\epsilon}\left(\frac{|1\rangle+|2\rangle}{\sqrt{2}}\otimes|0\rangle\right)=|B_{0}\rangle$. Note that the initial state is a product state while the final state has entanglement bounded away from zero, this transformation cannot be achieved by a unitary gate close to the identity. for a concrete example).
The above remark shows that an efficient local parameterization is not guaranteed even if the set of states is parameterized by only a constant number of parameters. As such, our goal – an efficient local parameterization of brickwork shallow-circuit states – is not as simple as it appears. We first use a lightcone argument to reduce the generating circuit into two effective layers, which is estimated up to a small error $\epsilon$. We then apply the inverse of the estimated outer layer gates to the state. Each gate in the outer layer of the generating circuit gets canceled out, up to an $O(\epsilon)$-rotation acting on a constant number of adjacent qubits. These $O(\epsilon)$-rotations are overlapping, but only with a constant of neighbors due to the locality of the shallow circuit. Therefore, the impact of each $O(\epsilon)$-rotation is limited to a constant-size neighborhood, and we can fully capture the impact by another $O(\epsilon)$-rotation with a constant number of parameters. Since there are only $O(n)$ such $O(\epsilon)$-rotations, putting these all together we get a parameter-efficient localization of the outer layer. A similar treatment can be applied to the inner layer. All together we get the desired result (see Appendix for details):
Theorem 1.
Let $\mathsf{S}_{{\rm sc}}$ be the subset of $n$-qubit states generated by depth-$d$ (1D) brickwork circuits. Given a fixed shallow-circuit state $|\psi_{0}\rangle=U_{0}|0\rangle^{\otimes n}$ of the same structure and the guarantee $d_{\operatorname{Tr}}(\psi_{{\rm sc}},\psi_{0})\leq\epsilon$ for every $|\psi_{{\rm sc}}\rangle\in\mathsf{S}_{{\rm sc}}$, there exists a set $\mathsf{W}_{\rm loc}$ of $O(\epsilon)$-rotations, parameterized by $O(n\cdot(2^{16d}/d))$ (real) local parameters, such that for every $|\psi_{{\rm sc}}\rangle\in\mathsf{S}_{\rm sc}$ there exists a $W\in\mathsf{W}_{\rm loc}$ with $|\psi_{\rm sc}\rangle=U_{0}W|0\rangle^{\otimes n}$. Here $W=\prod_{j=1}^{n_{\rm gate}}W^{28\epsilon}_{j}$ with $n_{\rm gate}\leq n/d+4$, and each $W^{28\epsilon}_{j}$ is a $(28\epsilon)$-rotation acting on no more than $8d$ neighboring qubits. On each qubit, no more than $n_{\rm overlap}=4$ gates in $\{W_{j}^{28\epsilon}\}$ act non-trivially.
Quantum local asymptotic normality (Q-LAN). Generally speaking, taking $N$ copies of a $D(<\infty)$-dimensional state as input, a Q-LAN [37, 38, 39] is an input-independent transformation that outputs a state close to a multi-mode Gaussian state. It is approximately reversible, as there exists an inverse Q-LAN recovering the original state up to an error vanishing in $N$. Here locality means that the Q-LAN works only if the concerned states are in a neighborhood with a vanishing radius. The processing of many-copy finite-dimensional quantum systems can then be reduced to that of Gaussian systems, which has been a key ingredient to achieve the ultimate precision limit of multiparameter quantum state estimation [40, 41, 42].
For our task, we would like to reduce $N$ copies of a locally-parameterized shallow-circuit state to a Gaussian state via the Q-LAN. A crucial issue with the existing Q-LANs is they require $D$ to be fixed with respect to $N$ ($n$), whereas we need $D$ to be growing exponentially fast in $n$. In addition, for existing Q-LANs, the parameterization of the $D$-dimensional state is set to a specific form (see [39]) that does not match ours. In the following, we bridge the gaps by proving a new Q-LAN.
We prove the result in a more general form that covers the reduced state in Theorem 1. First, define a circuit template to be $Q=(q_{1},\dots,q_{G})$, with each $q_{j}\subset[n]$ and $|q_{j}|\leq\tilde{d}$ for some $\tilde{d}$ independent of $n$. Let $\mathsf{S}_{N,\eta}(Q):=\left\{|\psi(\vec{H})\rangle^{\otimes N}~{}:~{}|\psi(\vec{H})\rangle=\prod_{j=1}^{G}e^{-i\frac{\eta}{\sqrt{N}}H_{j}}|0\rangle^{\otimes n}\right\}$ be the $N$-copy family of $n$-qubit pure states generated by $O(\eta/\sqrt{N})$-rotations with a fixed $Q$, where every $H_{j}$ is a Hermitian operator acting trivially on qubits not in $q_{j}$ and $\|H_{j}\|_{\infty}\leq 1$. Second, define $\mathsf{B}(Q)=\cup_{j=1}^{G}\left\{\vec{k}\in 2^{[n]}~{}:~{}k_{i}=0~{}\forall i\not\in q_{j}\right\}$ that keeps track of all $n$-qubit strings that could appear when acting $\sum_{j}H_{j}$ upon $|0\rangle^{\otimes n}$, and by
a coherent state with no more than $G\cdot 2^{\tilde{d}}$ modes, where $|u\rangle_{\rm coh}:=e^{-|u|^{2}/2}\sum_{m=0}^{\infty}\frac{u^{m}}{\sqrt{m!}}|m\rangle$ is a (single-mode) coherent state in an infinite-dimensional space spanned by the Fock basis $\{|m\rangle\}$, and $u^{\vec{k}}_{1}$ ($u^{\vec{k}}_{2}$) is the real (imaginary) part of $\eta\langle\vec{k}|\sum_{j=1}^{G}H_{j}|\vec{0}\rangle$. The amplitude of each mode of $|\vec{u}(\vec{H})\rangle_{\rm coh}$ is upper bounded by
$\max_{\vec{k}}|u^{\vec{k}}_{1}+iu^{\vec{k}}_{2}|\leq\eta\cdot n_{\rm overlap}$ (3)
where $n_{\rm overlap}:=\max_{i\in[n]}\sum_{j:i\in q_{j}}1$ is the maximum of the number of subsets overlapping at any local site. Our new Q-LAN is stated as follows:
Theorem 2.
When $\eta^{2}G^{2}\ll\sqrt{N}$, there exist quantum channels $\mathcal{V}_{N}$ (the Q-LAN) and $\mathcal{V}^{\ast}_{N}$ (the inverse Q-LAN) such that any $|\psi(\vec{H})\rangle\in\mathsf{S}_{N,\eta}(Q)$ can be reversibly converted into a coherent state (2) with no more than $G\cdot 2^{\tilde{d}}$ modes up to errors vanishing in $N$.
To establish the new Q-LAN, we first observe that the state can be expanded as $|\psi(\vec{H})\rangle=|0^{n}\rangle+\sum_{\vec{p}\in\mathsf{B}(Q)\setminus\{0^{n}\}}(\eta/\sqrt{N})a_{\vec{p}}|b\rangle+O(\eta^{2}/N)$ for some bounded $\{a_{\vec{p}}\}$. If $\eta^{2}/N$ is as small as assumed in Theorem 2, the difference between $|\psi(\vec{H})\rangle$ and $|\tilde{\psi}(\vec{H})\rangle\propto|0^{n}\rangle+\sum_{\vec{p}\in\mathsf{B}(Q)\setminus\{0^{n}\}}(\eta/\sqrt{N})a_{\vec{p}}|b\rangle$ will be negligible, so we can focus on constructing a Q-LAN for $|\tilde{\psi}(\vec{H})\rangle^{\otimes N}$. Note that $|\tilde{\psi}(\vec{H})\rangle$ lives in a much smaller space of dimension $|\mathsf{B}(Q)|=O(G)$. It follows that we may express $|\tilde{\psi}(\vec{H})\rangle^{\otimes N}$ in the symmetric basis, and define $\mathcal{V}_{N}$ to be one that maps symmetric states to multi-mode Fock states according to their types. More details can be found in Appendix.
Protocol 1 Compression protocol for $N$-copy shallow-circuit states.
1:Encoder:
2:Input: $N$ copies of a $n$-qubit state $|\psi_{\rm sc}\rangle$ generated by a brickwork circuit of depth-$d$;
3:a configuration of parameters $(\epsilon_{0},N_{0},\alpha_{0})$.
4:Require: A classical memory of $M_{\rm c}$ bits and a quantum memory of $M_{\rm q}$ qubits.
5:Construct an $\epsilon_{0}$-covering set $\mathsf{U}_{{\rm sc}}(\epsilon_{0})$ for brickwork circuits of depth-$d$. That is, for every $|\psi_{{\rm sc}}\rangle$, there exists an element of $\mathsf{U}_{{\rm sc}}(\epsilon_{0})$ whose trace distance with $|\psi_{\rm sc}\rangle$ is upper bounded by $\epsilon_{0}$.
6:Run state tomography with $N_{0}$ copies of $|\psi_{\rm sc}\rangle$, which outputs an estimate $|\hat{\psi}_{{\rm sc}}\rangle=\hat{U}|0\rangle^{\otimes n}$ with $\hat{U}\in\mathsf{U}_{{\rm sc}}(\epsilon_{0})$. Store $\hat{U}$ in the classical memory.
7:Apply first the inverse of $\hat{U}^{\otimes(N-N_{0})}$ and then the Q-LAN transformation on the remaining $N-N_{0}$ copies.
8:Amplify each mode of the resultant multi-mode coherent state: $|z\rangle\to|\sqrt{N/(N-N_{0})}z\rangle$, i.e., with intensity gain $N/(N-N_{0})$.
9:Compress the multi-mode coherent state via truncating each mode to less than $(e\alpha_{0})^{2}$ photons. Store the resultant state in the quantum memory.
1:Decoder:
5:Output the final state.
The compression protocol. With both tools in place, we are now ready to define our compression protocol, shown as Protocol 2.
In the End Matter, we give a high-level analysis of the protocol’s error and its memory cost, leaving details to Appendix. Essentially, we show that Protocol 2 with a configuration $N_{0}=\Theta(N^{1-\frac{1}{2}\Delta}),\epsilon_{0}=\Theta(n\cdot N^{-\frac{1}{2}(1-\frac{2}{3}\Delta)})$, and $\alpha_{0}=\Theta(\sqrt{N}\epsilon_{0})$ achieves both faithfulness and the desired compression rate, where $\Delta>0$ is a coefficient independent of $n$ and $N$.Another important question besides the memory efficiency is the computational efficiency. In Appendix, we show that Protocol 2 can be implemented within $\mathsf{poly}(n)$ time, via a high-dimensional Schur transform [43] and parallelizing the tomography [9] using the brickwork structure of generating circuits. It is a meaningful future direction of research to further reduce the complexity and to make the compression executable on practical devices.
Detailed error analysis (in the End Matter) also yields that $N$ must grow at least as a polynomial of $n$ with a high enough degree. Summarizing, we have the following theorem as the main result for compression:
Theorem 3.
Let $\mathsf{U}_{\rm sc}$ be the collection of all $n$-qubit pure states generated by depth-$d$ (1D) brickwork circuits, with $d$ being a fixed parameter. Assume $N=\Theta(n^{\frac{32}{3}+\gamma})$ for some $\gamma>0$. For any $\Delta\in\left(\frac{6}{32+3\gamma},\frac{3}{4}-\frac{18}{32+3\gamma}\right)$, there exist a time-efficient compression protocol that faithfully compresses $N$ copies of any state from $\mathsf{U}_{\rm sc}$ into a hybrid memory of $8(1-2\Delta/3)nd\left(\log_{2}N+O\left(\log_{2}n\right)\right)$ classical bits and $(n/d+4)2^{8d+1}/3\cdot\left(\Delta\log_{2}N+O(\log_{2}n)\right)$ qubits.
The total memory cost of our protocol is in the order of $n\cdot\log_{2}N$. One may ask whether the compression rate is optimal and if the cost can be made even smaller in terms of $n$ or $N$. Via an information-theoretic approach (details can be found in Appendix), we show that any faithful protocol requires a memory size close to the Holevo information [44] of the ensemble of shallow-circuit states, which is $\Omega(n\cdot\log_{2}N)$. This proves the optimality of the compression rate in the scaling of both $n$ and $N$.
The quantum-to-classical cost ratio (in the leading order) is $r_{\rm q-c}=\frac{2^{8d-2}\Delta}{d^{2}(3-2\Delta)}$ When the number of copies gets larger fixing $n$, i.e., when $\gamma\gg 1$, we may choose $\Delta\to 6/(32+3\gamma)$ and the ratio $r_{\rm q-c}=O(1/\gamma)$. This means that, although the sheer number of qubits in the hybrid memory might be large, it becomes negligible when compared to the number of bits. In this sense, when $N$ is large enough, $N$-copy shallow-circuit states can be converted into classical bits plus a small portion of qubits. The compressed $N$-copy state can be regarded as a long (classical) file with a quantum signature appended to it, similar as in the case of general qudits [33].
A reasonable suspicion then is whether any qubit is needed at all in the memory. A positive answer would render the compression much less interesting. Indeed, it would then be reduced to learning, where the encoder consists of tomography with all the $N$ copies and storing the outcome properly. Here we give a negative answer to this question and show the inability of any incoherent protocol. Explicitly, we show that any protocol using a fully classical memory cannot be faithful, no matter how large the classical memory is. This justifies the unique position of the task of compression as a fully coherent task, distinguishing it from the incoherent task of learning [9, 10, 11]. The proof idea Appendix, originally appeared in [33] to prove a similar result for qudits, is to keep track of genuinely quantum monotones such as the trace distance and the quantum Hellinger distance [45]. On one hand, we can find shallow-circuit states for which these monotones are strictly different, and the difference must be preserved by any faithful compression. On the other hand, these monotones coincide when the states are collapsed into bits, and the difference cannot be recovered, leading to a contradiction.
Applications and extensions. Our compression can store multiple instances of the same program at intermediate steps of a constant-depth computation or transfer them more efficiently in distributed settings. The need to compress $N$-copy states is also naturally encountered in new scenarios of quantum sensing, when the goal is to estimate the total effect of signals that occur in disjoint time intervals, e.g., the total duration of multiple spikes or the mean effect of intermittent fields. Our new protocol for shallow quantum circuits matches the trend to consider many-body probe states prepared by low-depth circuits [46]. The task is also a natural analogy of distributed quantum metrology [47, 48, 49, 50, 51, 52], where signals are now distributed in time instead of in space. There using an array of $N$ identical probes and the $N$-copy state compression to store the acquired information coherently and memory-efficiently in between time intervals achieves a substantial accuracy enhancement compared to measuring the signals separately, which has been verified in [32] for qubit probes used as a “quantum stopwatch”. Our protocol here enables multiparameter sensing using correlation-bounded many-body physical systems, which covers the detection of various electromagnetic properties. The application can even go beyond metrology: For example, the optimal tradeoff between the program size and the accuracy for programming a constant-size quantum gate is achieved via metrology [53], which might be extended to the programming of (1D brickwork) shallow quantum circuits using our compression as a subroutine.
@@
As we focused only on the most fundamental case, there is plenty of room for extension, e.g., to shallow-circuit states with a 2D structure. It is likely that we need to extend the tools, especially efficient local parameterization (Theorem 1), to 2D circuits. One possible route is to design new circuit covering schemes [54] for 2D or higher-dimension circuits that meet the requirement of local parameterization. Moreover, the circuit depth $d$ is treated as a constant throughout this work, but from the derivation of results it can be seen that the compression will still be faithful when $d$ grows very slowly (e.g., $d\ll\log n$) with $n$. In particular, it would be interesting to discuss pesudorandom quantum states [55], which are low-depth states processing approximate Haar-randomness and are thus of particular interest in quantum cryptography. At last, the extension to noisy setting is of particular interest in practice. While similar results are expected there, some techniques in this work do not immediately generalize to mixed states and require moderate adaptation.